# Tuning Algorithmic and Architectural Hyperparameters in Graph-Based Semi-Supervised Learning with Provable Guarantees

### v1

@@ 1 Introduction
Semi-supervised learning is a powerful paradigm in machine learning which reduces the dependence on expensive and hard-to-obtain labeled data, by using a combination of labeled and unlabeled data. This has become increasingly relevant in the era of large language models, where an extremely large amount of labeled training data is needed. A large number of techniques have been proposed in the literature to exploit the structure of unlabeled data, including popularly used graph-based semi-supervised learning algorithms (Blum and Mitchell, 1998; Zhu et al., 2003; Zhou et al., 2003; Delalleau et al., 2005; Chapelle et al., 2009). More recently, there has been an increasing interest in developing effective neural network architectures for graph-based learning (Kipf and Welling, 2017; Veličković, Petar et al., 2018; Iscen et al., 2019). However, different algorithms, architectures, and values of hyperparameters perform well on different datasets (Dwivedi et al., 2023), and there is no principled way of selecting the best approach for the data at hand. In this work, we study theoretically principled techniques for learning hyperparameters from infinitely large semi-supervised learning algorithm families.
In graph-based semi-supervised learning, the graph nodes consist of labeled and unlabeled data points, and the graph edges denote feature similarity between the nodes. Learning proceeds in two main stages—first a graph is designed by selecting which pairs of nodes have edges between them and how these edges are weighted, and next a graph partitioning algorithm is applied to label the unlabeled points while taking the graph structure into account. The second step is typically achieved by finding a partition that minimizes some objective function that typically involves hyperparameters. There are several classical ways of defining a graph-based regularization objective that depend on the available and predicted labels as well as the graph structure (Zhou et al., 2003; Delalleau et al., 2005). Optimizing this objective yields the predicted labels and the accuracy of the predictions depends on the chosen objective. The performance of the same objective may vary across datasets. By studying parameterized families of objectives, we can learn to design the objective that works best on a given domain-specific data. Similarly, modern deep learning based techniques often have several candidate architectures and choices for hyperparameters, often manually optimized for each application domain.
Recent work has considered the problem of learning the graph hyperparameter used in semi-supervised learning (i.e., the first stage of learning) (Balcan and Sharma, 2021; Fatemi et al., 2021) but leaves the problem of selecting the hyperparameter involved in the graph-based regularization objective that is used to actually label the points wide open. In this paper, we take important initial steps to build the theoretical foundations of algorithm hyperparameter selection in graph-based semi-supervised learning. Note that we focus specifically on labeling algorithm hyperparameters such as regularization strength, and architectural parameters such as self-loop weights. In contrast, optimization hyperparameters like learning rates and graph hyperparameters like kernel bandwidth are outside the scope of this study.
@@ 1.1 Contributions
  • We study hyperparameter tuning in three label propagation-based semi-supervised learning algorithms: the local and global consistency (Zhou et al., 2003), the smoothing-based (Delalleau et al., 2005), and a recently introduced normalized adjacency matrix based algorithm (Donnat and Jeong, 2023). We prove new $O\left(\log n\right)$ pseudo-dimension upper bounds for all three families, where $n$ is the number of graph nodes. Our proofs rely on a unified template based on determinant evaluation and root-counting, which may be of independent interest.
@@
  • We propose a new interpolated architecture, Graph Convolution and Attention Networks (GCAN), where a hyperparameter $\eta$ is introduced to interpolate two canonical GNN architectures: graph convolutional neural networks (GCNs) and graph attention networks (GATs). We bound the Rademacher complexity of tuning $\eta$. Our Rademacher complexity bounds for tuning SGC and GCAN have different dependencies on the feature dimension $d$: $\sqrt{d}$ for SGC, but $d$ for GCAN, corresponding to difference in the learning-theoretic complexities of the corresponding parameterized families.
  • We conduct experiments to demonstrate the effectiveness of our hyperparameter selection framework. Specifically, we show that GCAN can successfully compete with the better accuracy between GCN and GAT on each dataset, achieving the best-of-both-worlds consistently. We also evaluate the generalization error for tuning the label-propagation hyperparameters.
@@ 1.2 Related Work
Graph Based Semi-supervised Learning.
@@
Semi-supervised Learning is a popular machine learning paradigm with significant theoretical interest (Zhou et al., 2003; Delalleau et al., 2005; Garg et al., 2020). Classical algorithms focus on label-propagation based techniques, such as Zhou et al. (2003), Zhu et al. (2003), and many more. In recent years, graph neural networks (GNNs) have become increasingly popular in a wide range of application domains  (Kipf and Welling, 2017; Veličković, Petar et al., 2018; Iscen et al., 2019). A large number of different architectures have been proposed, including graph convolution networks, graph attention networks, message passing, and so on (Dwivedi et al., 2023). Both label propagation-based algorithms and neural network-based algorithms are useful in real life and perform well with appropriate hyperparameter tuning. For example, although GNN-based algorithms are more predominant in certain applications, Huang et al. (2021) show that modifications to label propagation-based algorithms can outperform GNN. For node classification in GNNs, many prior works study generalization guarantees for tuning network weights in GNNs  (Oono and Suzuki, 2020; Esser et al., 2021; Tang and Liu, 2023). In contrast, we study the tuning of the hyperparameters related to the GNN architecture.
Hyperparameter Selection.
@@
Hyper-parameters play an important role in the performance of both classical semi-supervised learning methods and GNNs. In general, hyperparameter tuning is performed on a validation dataset, and follows the same procedure: determine which hyperparameters to tune and then search within their domain for the combination of parameter values with best performance (Yu and Zhu, 2020). Many methods have been proposed to efficiently search within the parameter space, such as grid search, random search (Bergstra and Bengio, 2012), and Bayesian optimization (Mockus (1974); Mockus et al. (1978); Jones et al. (1998)). A few existing works investigate the theoretical aspects of these methods, such as through generalization guarantees and complexities of the algorithms. In particular, a recent line of work (Balcan et al., 2022, 2023) studies the tuning regularization hyperparameter in linear regression in statistical settings and provides generalization guarantees. For unsupervised learning, prior work (Balcan et al., 2018b, 2020a, 2024) has proposed several different parameterized algorithm families of clustering algorithms and studied the sample and computational complexity of learning the parameters. For semi-supervised learning, a recent line of work (Balcan and Sharma, 2021; Sharma and Jones, 2023) considers the problem of learning the best graph hyperparameter from a set of problem instances drawn from a data distribution. Another recent work (Balcan et al., 2025) investigates the kernel hyperparameters in GNN architectures, and derives the generalization guarantees through pseudo-dimension. However, no existing work theoretically studies the tuning of the graph partitioning (labeling) algorithm hyperparameter in semi-supervised learning, or investigates data-dependent bounds on hyperparameter selection in deep semi-supervised learning algorithms through Rademacher Complexity. We note that in this work we focus on the statistical learning setting (i.e. the problem instances are drawn from a fixed, unknown distribution), but it would be an interesting direction to study online tuning of the hyperparameters using tools from prior work (Balcan et al., 2018a, 2020b; Sharma and Suggala, 2025).
@@ 2 Preliminaries
Throughout this paper, $f(n)=O(g(n))$ denotes that there exists a constant $c>0$ such that $|f(n)|\leq c|g(n)|$ for all $n\geq n_{0}$. Similarly, $f(n)=\Omega(g(n))$ denotes that there exists a constant $c>0$ such that $|f(n)|\geq c|g(n)|$ for all sufficiently large $n$. The indicator function is indicated by $\mathbb{I}$, taking values in $\{0,1\}$. In addition, we define the shorthand $[c]=\{1,2,\ldots,c\}$. For a matrix $W$, we denote its Frobenius norm by $\|W\|_{F}$ and spectral norm by $\|W\|$. We also denote the Euclidean norm of a vector $v$ by $\|v\|$.
@@
An algorithm $F$ in this setting may be considered as a function that takes in $(X,Y)$ and outputs a predictor $f$ that predicts a label in $[c]$ for each data. We denote $f(z_{i})$ as our prediction on the $i$-th data. To evaluate the performance of a predictor $f$, we use 00-$1$ loss defined as $\frac{1}{n}\sum_{i=1}^{n}\ell_{0\text{-}1}\left(f(z_{i}),y_{i}\right)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}[f(z_{i})=y_{i}].$ In this work, we are interested in the generalizability w.r.t. the 00-$1$ loss.
@@
We consider several parameterized families of classification algorithms. Given a family of algorithms $\mathcal{F}_{\rho}$ parameterized by some hyperparameter $\rho\in\mathcal{P}$, and a set of $m$ problem instances $\{(X^{(k)},Y^{(k)})\}_{k=1}^{m}$ drawn i.i.d. from a fixed, unknown distribution $\mathcal{D}$ over datasets (with up to $n$ data points) on the input space $\mathcal{X}$ and the label space $\mathcal{Y}$, our goal is to select a parameter $\hat{\rho}$ such that the corresponding prediction function $f_{\hat{\rho}}$ of algorithm $F_{\hat{\rho}}$ minimizes the prediction error. That is, if we denote by $f_{\hat{\rho}}(z_{i}^{(k)})$ the predicted label for data point $z_{i}^{(k)}$ in the $k$-th problem instance, we want to study the generalization error of the hyperparameter $\hat{\rho}$ corresponding to Empirical Risk Minimization (ERM) over the problem instances (datasets) in our sample,
$\hat{\rho}=\text{argmin}_{\rho}\frac{1}{mn}\sum_{k=1}^{m}\sum_{i=1}^{n}\ell_{\mathrm{0\text{-}1}}(f_{\rho}(z_{i}^{(k)}),y_{i}^{(k)}).$
Each parameter value $\rho$ defines an algorithm $F_{\rho}$, mapping a problem instance $(X,Y)$ to a prediction function $f_{\rho}$, which induces a loss $H_{p}:=\frac{1}{n}\sum_{i=1}^{n}\ell_{0\text{-}1}(f_{\rho}(z_{i}),y_{i})$. Here $H_{\rho}$ is a function mapping a problem instance $(X,Y)$ to the corresponding average 0-1 loss. Define $\mathcal{H}_{\rho}=\{{H_{\rho^{\prime}}}\mid{\rho^{\prime}}\in\mathcal{P}\}$ as the class of loss functions parameterized by $\rho$.
Note that our problem setting differs from prior theoretical works on graph-based semi-supervised learning. The classical setting considers a single algorithm and learning the model parameter from a single problem instance. We are considering families of algorithms, each parameterized by a single hyperparameter, and aiming to learn the best hyperparameter across multiple problem instances. This is comparable to the setting introduced by Balcan and Sharma (2021), but they study tuning of the hyperparameters used in creating the graph instead. Our setting combines transductive and inductive aspects: each instance has a fixed graph of size $n$ (transductive), but the graphs themselves are drawn from an unknown meta-distribution (inductive).
@@
We study the generalization ability of several representative parameterized families of algorithms. That is, we aim to address the question of how many problem instances are required to learn a hyperparameter $\rho$ such that a learning algorithm can perform near-optimally for instances drawn from a fixed (but unknown) problem distribution. Clearly, the more complex the algorithm family, the more number of problem instances are needed.
Specifically, for each algorithm $f_{\hat{\rho}}$ trained given $m$ problem instances, we would like to bound the generalization error, i.e. the difference in the empirical 00-$1$ loss and the average 00-$1$ over the distribution:
$\frac{1}{mn}\sum_{k=1}^{m}\sum_{i=1}^{n}\ell_{\mathrm{0\text{-}1}}(f_{\hat{\rho}}(z_{i}^{(k)}),y_{i}^{(k)})-\mathbb{E}_{\left(X,Y\right)\sim\mathcal{D}}\left[\frac{1}{n}\sum_{i=1}^{n}\ell_{0\text{-}1}\left(f_{\hat{\rho}}(z_{i}),y_{i}\right)\right].$
To quantify this, we consider two learning-theoretic complexity measures for characterizing the learnability of algorithm families: the pseudo-dimension and the Rademacher complexity.
@@
Therefore, if we can show $\textsc{Pdim}(\mathcal{H}_{\rho})$ is bounded, then Theorem 2.1 implies using $m=O\left(\nicefrac{{\textsc{Pdim}(\mathcal{H})}}{{\epsilon^{2}}}\right)$ problem instances, the expected error on test instances is upper bounded by $\epsilon$ (with high probability). In Section 3, we will obtain optimal pseudo-dimension bounds for three different label-propagation algorithm families.
Another classical complexity measure is the Rademacher complexity, which is used to give data-dependent bounds:
@@
Given a space $\mathcal{X}$ and a sample $S=\{X^{(1)},\ldots,X^{(m)}\}\in\mathcal{X}^{m}$. Let $\mathcal{H}$ be the class of functions $H:\mathcal{X}\rightarrow\mathbb{R}$. The empirical Rademacher complexity of $\mathcal{H}$ is
@@
where $\sigma_{k}$ are Rademacher variables sampled uniformly and independently from $\{-1,1\}$.
@@
$\left|\frac{1}{m}\sum_{k=1}^{m}H({X^{(k)}})-\mathbb{E}_{X\sim D}[H(X)]\right|=O\left(\hat{R}_{m}(\mathcal{H})+\sqrt{\frac{1}{m}\log\frac{1}{\delta}}\right),~{}~{}~{}\text{for all}~{}~{}~{}H\in\mathcal{H}.$
To bound the Rademacher complexity in our setting, we restrict to binary classification $c=2$. For a predictor $f$, we also overload the notation and let $f(z_{i})\in[0,1]$ also be the output probability of node $z_{i}$ being classified as $1$. Instead of directly using the 00-$1$ loss, we upper bound it using margin loss, which is defined as
@@
Now we define $H^{\gamma}_{\rho}(X)=\frac{1}{n}\sum_{i=1}^{n}\ell_{\gamma}\left(f_{\rho}(z_{i}),y_{i}\right)$ to be the margin loss of the entire graph when using a parameterized algorithm $F_{\rho}$. Based on this definition, we have an induced loss function class $\mathcal{H}^{\gamma}_{\rho}$. Then, given $m$ instances, for any $\gamma>0$, we can obtain an upper bound for all $H_{\rho}^{\gamma}\in\mathcal{H}_{\rho}^{\gamma}$:
$\mathbb{E}_{\left(X,Y\right)\sim\mathcal{D}}$ $\left[\frac{1}{n}\sum_{i=1}^{n}\ell_{0\text{-}1}\left(f_{\hat{\rho}}(z_{i}),y_{i}\right)\right]$
@@
$=$ $\frac{1}{m}\sum_{i=1}^{m}H^{\gamma}_{\rho}(X^{(k)})+O\left(\hat{R}_{m}(\mathcal{H}_{\rho}^{\gamma})+\sqrt{\frac{\log{(1/\delta)}}{m}}\right)$ (by Theorem 2.2.)

### v2

@@ 1 Introduction
Semi-supervised learning is a powerful paradigm in machine learning which reduces the dependence on expensive and hard-to-obtain labeled data, by using a combination of labeled and unlabeled data. This has become increasingly relevant in the era of large language models, where an extremely large amount of labeled training data is needed. A large number of techniques have been proposed in the literature to exploit the structure of unlabeled data, including popularly used graph-based semi-supervised learning algorithms (Blum and Chawla, 2001; Zhu et al., 2003; Zhou et al., 2003; Delalleau et al., 2005; Chapelle et al., 2009). More recently, there has been an increasing interest in developing effective neural network architectures for graph-based learning (Kipf and Welling, 2017; Veličković, Petar et al., 2018; Iscen et al., 2019). However, different algorithms, architectures, and values of hyperparameters perform well on different datasets (Dwivedi et al., 2023), and there is no principled way of selecting the best approach for the data at hand. In this work, we initiate the study of theoretically principled techniques for learning hyperparameters from infinitely large semi-supervised learning algorithm families.
In graph-based semi-supervised learning, the graph nodes consist of labeled and unlabeled data points, and the graph edges denote feature similarity between the nodes. There are several classical ways of defining a graph-based regularization objective that depend on the available and predicted labels as well as the graph structure. Optimizing this objective yields the predicted labels and the accuracy of the predictions depends on the chosen objective. The performance of the same objective may vary across datasets. By studying parameterized families of objectives, we can learn to design the objective that works best on a given domain-specific data. Similarly, modern deep learning based techniques often have several candidate architectures and choices for hyperparameters, often manually optimized for each application domain. Recent work has considered the problem of learning the graph hyperparameter used in semi-supervised learning (Balcan and Sharma, 2021; Fatemi et al., 2021) but leaves the problem of selecting the hyperparameter wide open. In this paper, we take important initial steps to build the theoretical foundations of algorithm hyperparameter selection in graph-based semi-supervised learning.
Note that we focus specifically on algorithm hyperparameters, such as self-loop weights, leaving optimization hyperparameters like learning rates outside the scope of this study.
@@ 1.1 Contributions
  • We study hyperparameter tuning in three canonical label propagation-based semi-supervised learning algorithms: the local and global consistency (Zhou et al., 2003), the smoothing-based (Delalleau et al., 2005), and a novel normalized adjacency matrix-based algorithm. We prove new $O\left(\log n\right)$ pseudo-dimension upper bounds for all three families, where $n$ is the number of graph nodes. Our proofs rely on a unified template based on determinant evaluation and root-counting, which may be of independent interest.
@@
  • We propose an architecture (GCAN) where a hyperparameter $\eta$ is introduced to interpolate two canonical GNN architectures: graph convolutional neural networks (GCNs) and graph attention neural networks (GATs). We bound the Rademacher complexity of tuning $\eta$. Because the parameter dimension is different, the Rademacher complexity of SGC and GCAN has different dependencies on the feature dimension $d$: $\sqrt{d}$ for SGC while $d$ for GCAN.
  • We conducted experiments to demonstrate the effectiveness of our hyperparameter selection framework.
@@ 1.2 Related Work
Graph Based Semi-supervised Learning
@@
Semi-supervised learning is a popular machine learning paradigm with significant theoretical interest (Zhou et al., 2003; Delalleau et al., 2005; Balcan and Blum, 2010; Garg et al., 2020). Classical algorithms focus on label-propagation based techniques, such as Zhou et al. (2003), Zhu et al. (2003), and many more. In recent years, graph neural networks (GNNs) have become increasingly popular in a wide range of application domains  (Kipf and Welling, 2017; Veličković, Petar et al., 2018; Iscen et al., 2019). A large number of different architectures have been proposed, including graph convolution networks, graph attention networks, message passing, and so on (Dwivedi et al., 2023). Both label propagation-based algorithms and neural network-based algorithms are practically useful (Balcan et al., 2005; Kipf and Welling, 2017). For example, although GNN-based algorithms are more predominant in applications, Huang et al. (2020) show that modifications to label propagation-based algorithms can outperform GNN. For node classification in GNN, many work study generalization guarantees for tuning network weights in GNNs  (Oono and Suzuki, 2021; Esser et al., 2021; Tang and Liu, 2023). In contrast, we study the tuning of the hyperparameters related to the GNN architecture.
Hyperparameter Selection
@@
Hyper-parameters, such as the weight for self-loop, play important roles in the performance of both classical methods and GNNs. In general, hyperparameter tuning is performed on a validation dataset, and follows the same procedure: determine which hyperparameters to tune and then search within their domain for the combination of parameter values with best performance (Yu and Zhu, 2020). Many methods are proposed to efficiently search within the parameter space, such as grid search, random search (Bergstra and Bengio, 2012), and Bayesian optimization (Mockus (1974); Mockus et al. (1978); Jones et al. (1998)). A few existing works investigate the theoretical aspects of these methods, such as through generalization guarantees and complexities of the algorithms.
A recently introduced paradigm called data-driven algorithm design is useful for obtaining formal guarantees for hyperparameter tuning (Balcan, 2020; Sharma, 2024b), and has found several applications (for instance, Blum et al. 2021; Bartlett et al. 2022; Balcan et al. 2023a, 2024). In particular, Balcan et al. (2022, 2023b, 2025a) study the regularization hyperparameter in the ElasticNet and Balcan and Sharma (2024) study learning decision tree algorithms. For unsupervised learning, Balcan et al. (2018, 2025b) study a parameterized family of clustering algorithms and study the sample and computational complexity of learning the parameters. For semi-supervised learning, a recent line of work (Balcan and Sharma (2021); Sharma and Jones (2023)) considers the problem of learning the best graph hyperparameter from a set of problem instances drawn from a data distribution. Another recent work (Balcan et al., 2025c) investigates the kernel hyperparameters in GNN architecture, and derives the generalization guarantees through pseudo-dimension. However, no existing work theoretically studies the tuning of the labeling algorithm hyperparameter in semi-supervised learning, or investigates data-dependent bounds on hyperparameter selection in deep semi-supervised learning algorithms through Rademacher Complexity. We note that in this work we focus on the statistical learning setting (i.e. the problem instances are drawn from a fixed, unknown distribution), but it would be an interesting direction to study online tuning of the hyperparameters using tools from prior work (Balcan et al., 2019; Sharma, 2024a; Sharma and Suggala, 2025).
@@ 2 Preliminaries
Throughout this paper, $f(n)=O(g(n))$ denotes that there exists a constant $c>0$ such that $|f(n)|\leq c|g(n)|$. $f(n)=\Omega(g(n))$ denotes that there exists a constant $c>0$ such that $|f(n)|\geq c|g(n)|$. The indicator function is indicated by $\mathbb{I}$, taking values in $\{0,1\}$. In addition, we define the shorthand $[c]=\{1,2,\ldots,c\}$. For a matrix $W$, we denote its Frobenius norm by $\|W\|_{F}$ and spectral norm by $\|W\|$. We also denote the Euclidean norm of a vector $v$ by $\|v\|$.
@@
An algorithm $F$ in this setting may be considered as a function that takes in $(X,Y)$ and outputs a predictor $f$ that predicts a label in $[c]$ for each data. We denote $f(z_{i})$ as our prediction on the $i$-th data. To evaluate the performance of a predictor $f$, we use 00-$1$ loss (i.e. the predictive inaccuracy) defined as $\frac{1}{n}\sum_{i=1}^{n}\ell_{0-1}\left(f(z_{i}),y_{i}\right)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}[f(z_{i})\neq y_{i}].$ In this work, we are interested in the generalizability of an algorithm $F$ on 00-$1$ loss.
@@
We consider several parameterized families of classification algorithms. Given a family of algorithms $\mathcal{F}_{\rho}$ parameterized by some parameter $\rho$, and a set of $m$ problem instances $\{(X^{(k)},Y^{(k)})\}_{k=1}^{m}$ i.i.d. generated from the data distribution $\mathcal{D}$ of the input space $\mathcal{X}$ and the label space $\mathcal{Y}$, our goal is to select a parameter $\hat{\rho}$ whose corresponding prediction function $f_{\hat{\rho}}$ of algorithm $F_{\hat{\rho}}$ minimizes the prediction error. That is, denote $f_{\hat{\rho}}(z_{i}^{(k)})$ as the predicted label of data point $z_{i}^{(k)}$ in the $k$-th problem instance, we want
$\hat{\rho}=\text{argmin}_{\rho}\frac{1}{mn}\sum_{k=1}^{m}\sum_{i=1}^{n}\ell_{\mathrm{0-1}}(f_{\rho}(z_{i}^{(k)}),y_{i}^{(k)}).$
Each parameter value $\rho$ defines an algorithm $F_{\rho}$, mapping a problem instance $(X,Y)$ to a prediction function $f_{\rho}$, which induces a loss $\frac{1}{n}\sum_{i=1}^{n}\ell_{0-1}(f_{\rho}(z_{i}),y_{i})$. We define $H_{\rho}$ as the function mapping $(X,Y)$ to this loss and $\mathcal{H}\rho={H{\rho^{\prime}}}_{\rho^{\prime}}$ as the family of loss functions parameterized by $\rho$.
Note that our problem setting differs from prior theoretical works on graph-based semi-supervised learning. The classical setting considers a single algorithm and learning the model parameter from a single problem instance. We are considering families of algorithms, each parameterized by a single hyperparameter, and aiming to learn the best hyperparameter across multiple problem instances. Our setting combines transductive and inductive aspects: each instance has a fixed graph of size $n$ (transductive), but the graphs themselves are drawn from an unknown meta-distribution (inductive).
@@
We study the generalization ability of several representative parameterized families of algorithms. That is, we aim to address the question of how many problem instances are required to learn a hyperparameter $\rho$ such that a learning algorithm can perform near-optimally for instances drawn from a fixed problem distribution. Clearly, the more complex the algorithm family, the more number of problem instances are needed.
Specifically, for each algorithm $f_{\hat{\rho}}$ trained given $m$ problem instances, we study the difference in the empirical 00-$1$ loss and the actual 00-$1$ on the distribution:
$\frac{1}{mn}\sum_{k=1}^{m}$ $\sum_{i=1}^{n}\ell_{\mathrm{0-1}}(f_{\rho}(z_{i}^{(k)}),y_{i}^{(k)})$
-- $\mathbb{E}_{\left(X,Y\right)\sim\mathcal{D}}\left[\frac{1}{n}\sum_{i=1}^{n}\ell_{0-1}\left(f_{\hat{\rho}}(z_{i}),y_{i}\right)\right].$
To quantify this, we consider two complexity measures for characterizing the learnability of algorithm families: the pseudo-dimension and the Rademacher complexity.
@@
Therefore, if we can show $\textsc{Pdim}(\mathcal{H}_{\rho})$ is bounded, then using the standard empirical risk minimization argument, Theorem 2.1 implies using $m=O\left(\nicefrac{{\textsc{Pdim}(\mathcal{H})}}{{\epsilon^{2}}}\right)$ problem instances, the expected error on test instances is upper bounded by $\epsilon$. In Section 3, we will obtain optimal pseudo-dimension bounds for three canonical label-propagation algorithm families.
Another classical complexity measure is the Rademacher complexity:
@@
Given a space $\mathcal{X}$ and a distribution $\mathcal{D}$, let $S=\{X^{(1)},\ldots,X^{(m)}\}$ be a set of examples drawn i.i.d. from $\mathcal{D}$. Let $\mathcal{H}$ be the class of functions $H:\mathcal{X}\rightarrow\mathbb{R}$. The (empirical) Rademacher complexity of $\mathcal{H}$ is
@@
where each $\sigma_{k}$ is i.i.d. sampled from $\{-1,1\}$.
@@
$\left|\frac{1}{m}\sum_{k=1}^{m}H({X^{(k)}})-\mathbb{E}_{X\sim D}[H(X)]\right|$
$\qquad=O\left(\hat{R}_{m}(\mathcal{H})+\sqrt{\frac{1}{m}\log\frac{1}{\delta}}\right),~{}~{}~{}\text{for all}~{}~{}~{}H\in\mathcal{H}.$
To bound the Rademacher complexity in our setting, we restrict to binary classification $c=2$ and change the label space to $Y\in\{-1,1\}^{n}$. For a predictor $f$, we also overload notation and let $f(z_{i})\in[0,1]$ be the output probability of node $z_{i}$ being classified as $1$. Instead of directly using the 00-$1$ loss, we upper bound it using margin loss, which is defined as
@@
Now we define $H^{\gamma}_{\rho}(X)=\frac{1}{n}\sum_{i=1}^{n}\ell_{\gamma}\left(f_{\rho}(z_{i}),y_{i}\right)$ to be the margin loss of the entire graph when using a parameterized algorithm $F_{\rho}$. Based on this definition, we have an induced loss function family $\mathcal{H}^{\gamma}_{\rho}$. Then, given $m$ instances, for any $\gamma>0$, we can obtain an upper bound for all $H_{\rho}^{\gamma}\in\mathcal{H}_{\rho}^{\gamma}$:
$\mathbb{E}$ [${}_{\left(X,Y\right)\sim\mathcal{D}}\left[\frac{1}{n}\sum_{i=1}^{n}\ell_{0-1}\left(f_{\hat{\rho}}(z_{i}),y_{i}\right)\right]$
@@
$=\frac{1}{m}\sum_{i=1}^{m}H^{\gamma}_{\rho}(X^{(k)})+O\left(\hat{R}_{m}(\mathcal{H}_{\rho}^{\gamma})+\sqrt{\frac{\log{(1/\delta)}}{m}}\right).$ (by Theorem 2.2)
@@
Note that these guarantees we obtained can also be applied to some standard hyperparameter tuning methods like grid search with cross-validation. For example, in k-fold cross-validation, if we define the distribution of problem instances as a uniform distribution on these small subsets of validation data, then our results imply the necessary number of iterations $k$ needed to effectively tune hyperparameters using cross-validation. That is, the number of folds of cross-validation needed to learn a hyperparameter that performs nearly as well as the hyperparameter if cross-validation were run to convergence.